A Bibliography of Books and Manuals on the Oracle Database System

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

12 August 2014 Version 2.03

Abstract

This bibliography records books about the Oracle database system.

Title word cross-reference

*Forms [McC94].

3 [HB98].

7 [ALZ94, Cha96, CCD^{+}95, Jai97, Lue95, Ora95, Ora97b, Ora97c, Ora97a, Ora97g, Stu95a, Stu95b]. 7.0 [Aul94]. 7.3 [S^{+}97].

8 [Tja97].

95 [Jai97].

Administer [Ora97b, Ora97c, Ora97a]. Administration [ALS97, Bur99, KL99, ALZ94, Aul94, Aul97, Aul98b, Bro97, Bro99, Cha96]. administrator [Lon94b]. Advanced [ALS97, Feu96, CLC^{+}95]. analysis [CA97a]. Application [BNT98, Dyn98, HB98]. Applications [Mue00, Ode99, Bur96, Hip96]. approach [BB93]. ARC [Gua97]. ARC-to-Oracle [Gua97]. architecture [ALZ94, Bob98]. automation [Ing97].

Backup [RC04, Vel97, VA98, A^{+}99, Vel95]. Beginner [AC95, CA97b, AC97]. Best [Bon99, Feu01]. between [Wil97]. black [Aul98c]. book [Aul98c]. Bookshelf [Ass00]. build [CA97a]. Building [Mue00, Owe96]. Built [FDB98, FBD98]. Built-in [FDB98]. Built-ins [FBD98].

C [WL91]. CASE [BB93, Smi94]. CD [Ass00]. Checklists [Rev01]. client [Hip96]. client/server [Hip96]. Cluster [Don02]. Code [Feu01]. Complete [KL97a, KL97b, KL95]. Concevoir [Smi94]. Configuration [TG00]. construction [Pal97]. cooperative [Kli95]. Corporation [Wil97]. cours [MPK97]. cram [Aul98a]. Create [PB97]. Critical [Bon99].

Data [CA97a, CAAT98]. Database [KL99, PB97, Rod99, Stu97, Bob98, Bro97, Bro99, Bur96, Ing97, Loc97, Lon94b, Ora97b, Ora97c, Ora97a, Pal97]. Databases [PB97, RS99, Owe96]. days [Loc97, WD98]. DBA [Aul98a, A^{+}99, Gre95, Kre05, Lon94b, Lon94a, Lon98, Lonxx, MM98, Rev01, RC04]. DBAs [Bur01, DS02]. dBase [Car93]. Definitive [Gen99, GM01]. Design [ES97a, ES97b, Tja97]. Designer [AW97, Cob95, DK97]. Designer/2000 [AW97, Cob95, DK97]. Detailed [CCD^{+}95]. Developer [FO00, FSG97, HB98, Ode99, Ora97f, Rod99, Ora94, CL95, H^{+}95, HH91, McC96, S^{+}97, Stu95a, Stu95b, Cob95, CP95, Hip96, MPK97, MM98, Mul96, Mul97, Ora97d, Ora97e, Sto96]. Developer/2000 [Ora97f, Cob95, CP95, Hip96, MPK97, MM98, Mul96, Mul97, Ora97d, Ora97e, Sto96]. Developing [Hip96]. development [BB93, Kli95, Lak02, Loc97, WL91]. developper [Smi94]. difference [Wil97]. Distributed [Dye99, WL91].

Edition [KL97a, Vel97, Lon94a]. effective [Lon94b]. efficient [Lon94b]. Electronic [KL97a]. Ellison [Wil97]. Engineering [Bon99]. Engines [PB97]. environment [Kli95]. Essential [KL99, Kre00, Lue95, Urm97]. Essentials [GSS99, GSS01]. Evaluating [Gua97]. every [Lon94b, Urm97]. exam [Aul98a]. Example [RS04]. Experiences [RS99]. Exploring [Cob95]. express [Tja97]. Extensible [RS99].

Features [Feu99]. formation [Cha96].

Forms [CLC^{+}95, CL95, Ora97e].

Fundamentals [Pat99].

generation [CP95]. God [Wil97]. Graphics [H^{+}95, Sel95]. Guide [AC95, Cha96, CA97b, Feu99, FSG97, Gen99, GM01, Gen02, HB98, Rod99, AC97, CL95, CP95, CA97a, Gre95, HM92, H^{+}95, HH91, Kli95, Kre05, McC96, MM98, S^{+}97, Stu95a, Stu95b, Urm97, WL91].

Handbook

[Dyn98, Lonxx, Mul97, Vel97, VA98, AW97, Cha00, CSK01, Cox95, DK97, KP95, Lon94b, Lon94a, Lon98, Mul96, Sto96, Vel95]. High [Bur96, Ing97]. High-performance [Bur96].

II [Ora97b, Ora97c]. implementation [CCD^{+}95, Tja97]. including [MM98, Stu95a, Stu95b]. information [Stu97, Urm97]. Ingres [Car93]. ins [FBD98]. intelligent [Owe96]. Interface [PB97, SAS93]. Internal [Ada99]. Introduction [Ode99, McC94, Ora97e]. IV [Car93].

Java [Bal02, Lak02, Pri01]. JDBC [Bal02].

kit [Pal97]. know [Lon94b].

langages [ALZ94]. Language [FPD99]. Larry [Wil97]. Latches [Ada99]. Learn [PB97]. Learning [PF02]. linkages [Gua97]. Locks [Ada99].

management [Aul94, Aul97, Aul98b, Lon94b, Sob97]. manual [Sel95]. mapping [Tja97]. Mastering [GH97, McD03, MB02]. Memory [Ada99]. migrate [Bob98]. Mission [Bon99]. Mission-Critical [Bon99].

needs [Lon94b]. Net8 [TG00]. Networking [Tol96]. NIST [Sob97]. NT [Bro99]. Nutshell [GK02].

Objects [FSG97, GH97, KP95]. Open [DH01]. Optimizing [Feu01]. Oracle [CAD95, HH87, Owe96, Urm97, CG93, HM92, SAS93, Sob97, AC95, ALZ94, Alo97, AW97, Ora94, ALS97, Aul94, Aul98a, A^{+}99, Bal02, BB93, Bob98, Bon99, Bro97, BNT98, Bur96, Bur99, Bur01, Car93, Cha00, Cha96, CCD^{+}95, Cob95, CA97a, Cox95, Dyn98, Del94, DK97, DH01, DS02, Dye99, ES97a, Feu95, Feu96, FP97, FDB98, FBD98, FPD99, Feu99, FO00, Feu01, FP02, FP05, FSG97, Gen99, Gen00, GM01, Gen02, Gre95, GH97, GSS99, GSS01, GK02, Gua97, GC96, Gur02, HB98, Hip96, HH89, HH91, Ing97, Jai97, Kli95, KL95, KL97a, KP95, KL99, Kre00, Kre05, Kro93, KS02, Lak02, LK98, Lon94b, Lon94a, Lue95, MM00, McC96, McC94, MM02, MSMS05, McD03, MB02, MPK97, MM98, Mue00, Mul96, Mul97, Ass00]. Oracle [Ode99, Pal97, PB97, PF02, Pri01, Rev01, Rod99, RS04, RC04, S^{+}97, Smi94, Sto96, Stu95a, Stu95b, TH98, Tja97, Tol96, TG00, Urm96, Vel95, Vel97, VA97, WL91, Wil97]. ORACLE/SQL [HM92]. Oracle8 [AC97, Alo98, Aul97, Aul98b, Aul98c, Bob98, Bro99, CA97b, CAAT98, CADA98, ES97b, GSS99, GSS01, KL97b, Loc97, Lon98, Lonxx, PH98, Urm97, VA98, WD98]. Oracle8i [Ada99, Alo01, Feu99, GSS99, RS99]. Oracle9 [GSS01]. Oracle9i [Don02, CSK01].

Packages [Feu96, FDB98]. Pages [PB97]. Paradox [Jai97]. Parallel [MM00]. 2000 [Ora97f, AW97, Cob95, CP95, DK97, Hip96, MPK97, MM98, Mul96, Mul97, Ora97d, Ora97e, Sto96]. ACCESS [SAS93]. server [Hip96]. SQL [Ass00, Feu95, Feu96, FP97, FBD98, FPD99, Feu99, FO00, Feu01, FP02, HM92, McC94, McD03, Ode99, Owe96, PF02, RS04, Urm96, Urm97]. Performance [Alo01, CG93, GC96, Alo97, Alo98, A^{+}99, Bur96, Ing97, Urm97]. Perl [DS02]. Personal [Ora95, Jai97]. PL [Ass00, Feu95, Feu96, FP97, FBD98, FPD99, Feu99, FO00, Feu01, FP02, FP05, McC94, McD03, Ode99, Owe96, PF02, RS04, Urm96, Urm97]. PL/ SQL [Ass00, Feu95, Feu96, FP97, FBD98, FPD99, Feu99, FO00, Feu01, FP02, McC94, McD03, Ode99, Owe96, PF02, RS04, Urm96, Urm97]. PL/SQL [FP05]. plan [Bob98, CCD^{+}95]. Pocket [FBD98, FPD99, Gen00, Gen02, Gur02, KS02, MM02, Rev01, Bur01, Kre05, MSMS05]. Power [FSG97, GH97, KP95]. Practical [McD03, CA97a]. Practices [Bon99, Feu01]. pratiques [MPK97]. procedure [Owe96]. Processing [MM00]. professional [HM92]. Programmer [Del94, HM92, Urm97, WL91]. Programming [Bal02, Car93, Feu95, Feu96, FP97, Feu99, FP02, FP05, Pri01, Urm96, Urm97].

Quick [RC04, Don02].

Rapid [BB93]. Reader [Stu97]. Recovery [RC04, Vel97, VA98, A^{+}99, Vel95]. Reference [BNT98, FBD98, FPD99, Gen00, Gen02, Gur02, KL97a, KL97b, KL99, Kre00, KS02, MM02, Rev01, RC04, Bur01, Don02, Kli95, KL95, MSMS05, SAS93, Sel95]. release [CLC^{+}95, CL95, CP95, H^{+}95, Sel95, Stu95a]. reports [Ora97f]. retrieval [Stu97]. revolutionary [Bob98]. RMAN [KS02]. roll [CA97a]. roll-out [CA97a].

SAP [Bur99]. SAS/ACCESS [SAS93]. Scripts [LK98]. Security [TH98]. Server [BNT98, Dyn98, HB98, Cox95, Ora97g]. Services [Ada99]. Software [Bon99]. Solutions [McD03]. Source [DH01]. space [Sob97]. Spatial [RS99]. SQL [Car93, FP05, Gur02, HH91, Kre00, McC94, MB02, Pat99]. SQL*Forms [McC94]. SQL*Loader [GM01]. SQL*Plus [Gen99, Gen00, Gen02]. SQLJ [Pri01]. stored [Owe96]. successful [CA97a]. survival [Gre95]. Syntax [Gen02]. system [Lon94b, Sob97]. Systems [Bon99, Dye99, Bro97, WL91].

Teach [Loc97, WD98]. techniques [CLC^{+}95]. test [A^{+}99]. those [McC94]. Tips [ES97b]. TOAD [MM02, MSMS05]. Toolkit [BNT98]. toolset [Kro93]. travaux [MPK97]. triggers [Owe96]. Troubleshooting [TG00, VA97]. Tru64 [Don02]. Tuning [Alo01, ALS97, CAD95, CADA98, CG93, GC96, Gur02, Alo97, Alo98, Urm97]. turning [A^{+}99].

Understand [Bob98]. UNIX [Don02, Alo97, Alo98, Alo01, Bro97, Bur01]. usage [SAS93]. user [Kli95, Stu95a, Stu95b]. Using [Kro93, PH98, McC94, Ora97f, Sob97]. utilisation [Cha96]. utilities [MM98].

version [HH89, McC94, SAS93, Stu95b]. versions [Del94, Smi94]. vs [Jai97].

Waits [Ada99]. warehouse [CA97a]. Warehousing [CA97a, CAAT98]. Web [HB98, BNT98, Dyn98, Ode99, PB97, Sob97]. WebServer [MM98]. Wide [Sob97]. Windows [Bro99, Jai97]. Workbook [FO00]. workgroup [Cox95, Ora97g]. Working [HH87, HH89]. workshop [BB93].

World [Sob97].

XML [Cha00, CSK01, Mue00].

yourself [Loc97, WD98].

References

Ault:1999:ODT

[A^{+}99] Michael R. Ault et al., editors. Oracle DBA: test 3 and test 4: backup and recovery, performance turning. Coriolis Group Books, Scottsdale, AZ, USA, 1999. ISBN 1-57610-331-5. xxiii + 385 pp. LCCN QA76.9.D3 O714 1999.

Abbey:1995:OBG

[AC95] Michael Abbey and Michael J. Corey. Oracle: A Beginner’s Guide. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 1995. ISBN 0-07-882122-3. 552 pp. LCCN QA76.9.D3A25 1995. US$29.95. URL http://

                                                www.osborne.com/oracle/index.

htm. Foreword by Gary E. Damiano.

Abbey:1997:OBG

[AC97] Michael Abbey and Michael J. Corey. Oracle8: a beginner’s guide. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 1997. ISBN 0-07-882393-5. xxviii + 767 pp. LCCN QA76.9.D3 A24

                                                1997. URL http://www.osborne.
                                                com/oracle/index.htm.

Adams:1999:OIS

[Ada99] Steve Adams. Oracle8i Internal Services for Waits, Latches, Locks, and Memory. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-598-X. 120 pp. US$19.95.

URL http://www.oreilly.com/
catalog/orinternals.

Alomari:1997:OUP

[Alo97] Ahmed Alomari. Oracle and UNIX performance tuning. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-849167-4. xxiii + 255 pp. LCCN QA76.9.D3 A52 1997.

Alomari:1998:OUP

[Alo98] Ahmed Alomari. Oracle8 and UNIX performance tuning. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1998. ISBN 0-13-907676-X. ???? pp. LCCN QA76.9.D3 A519 1998. US$39.95.

URL http://www.prenhall.com/
allbooks/ptr_013907676x.html.

Alomari:2001:OUP

[Alo01] Ahmed Alomari. Oracle8i and UNIX Performance Tuning. Prentice Hall PTR Oracle series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2001. ISBN 0-13-018706-2. xxix + 415 pp. LCCN QA76.9.D3 A519 2001. US$49.99. URL http:

//www.phptr.com/ptrbooks/ptr_
0130187062.html.

Aronoff:1997:AOT

[ALS97] Eyal Aronoff, Kevin Loney, and Noorali Sonawalla. Advanced Oracle Tuning and Administration. Osborne/McGraw-Hill, Berkeley, CA, USA, 1997. ISBN 0-07-882241-6. xxvi + 595 pp. LCCN QA76.9.D3A735 1997. US$39.99.

                                        URL http://www.osborne.com/
                                        oracle/index.htm.

Abdellatif:1994:OLA

[ALZ94] Abdelaziz Abdellatif, Mohamed Limame, and Abdelmalek Zeroual. Oracle 7: langages, architecture, administration. Eyrolles, Paris, France, 1994. ISBN 2-212-08832-9. xi + 464 pp. LCCN ????

OReilly:2000:OPS

[Ass00] O’Reilly and Inc. Associates. The Oracle PL/SQL CD Bookshelf. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, July 2000. ISBN 1-56592-849-0. 285 pp. LCCN ???? US$89.95.

                                        URL http://www.oreilly.com/
                                        catalog/oraclecdbs. Includes CD-

ROM.

Ault:1994:OAM

[Aul94] Michael R. Ault. Oracle 7.0 administration & management. Wiley, New York, NY, USA, 1994. ISBN 0-471-60857-2 (paperback). xxi + 516 pp. LCCN QA76.9.D3 A937 1994.

Ault:1997:OAM

[Aul97] Michael R. Ault. Oracle8 administration and management. Wiley, New York, NY, USA, 1997. ISBN 0-471-19234-1 (paperback). ???? pp. LCCN QA76.9.D3 A938 1997.

Ault:1998:ODE

[Aul98a] Michael R. Ault. Oracle DBA exam cram. Coriolis Group Books, Scottsdale, AZ, USA, 1998. ISBN 1-57610-262-9. xviii + 413 pp. LCCN QA76.9.D3 A9395 1998.

Ault:1998:OAM

[Aul98b] Michael R. Ault. Oracle8 administration and management. Wiley Computer Publishers, New York, NY, USA, 1998. ISBN 0-471-19234-1 (paperback ). xli + 836 pp. LCCN QA76.9.D3A938 1997.

Ault:1998:OBB

[Aul98c] Michael R. Ault. Oracle8 black book.

Coriolis Group Books, Scottsdale, AZ, USA, 1998. ISBN 1-57610-187-8. xix + 906 pp. LCCN QA76.9.D3 A939 1998.

Anderson:1997:ODH

[AW97] Carrie Anderson and David Wendelken. The Oracle Designer/2000 handbook. Addison-Wesley, Reading, MA, USA, 1997. ISBN 0-201-63445-7. xviii + 606 pp. LCCN QA76.758.A55 1997.

Bales:2002:JPO

[Bal02] Donald Bales. Java Programming with Oracle JDBC. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00088-X. xv + 477 pp. LCCN QA76.73.J38 B35 2002 Stacks. US$39.95.

URL  http://safari.oreilly.

com/059600088X;   http://www.
oreilly.com/catalog/jorajdbc.

Billings:1993:RDO

[BB93] Chris Billings and Maria Billings. Rapid development with Oracle CASE: a workshop approach. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-63344-2. xii + 129 pp. LCCN QA76.76.D47B55 1993.

Brown:1998:OAS

[BNT98] Bradley D. Brown, Richard J. Niemiec, and Joseph C. Trezzo. Oracle Application Server Web Toolkit Reference. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 1998. ISBN 0-07-882433-8. xxxv + 924 pp. LCCN QA76.9.D3B774 1998. US$44.99.

URL http://www.osborne.com/
oracle/index.htm.

Bobrowski:1998:OAU

[Bob98] Steven M. Bobrowski. Oracle8 architecture: Understand, plan, and migrate to Oracle’s revolutionary database. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 1998. ISBN 0-07-882274-2 (paperback). xxiv + 356 pp. LCCN QA76.9.D3B6193 1998. US$34.99. URL http://

www.osborne.com/oracle/index.

htm. Foreword by Ken Jacobs.

Bonazzi:1999:SEO

[Bon99] Elio Bonazzi. Software Engineering with Oracle: Best Practices for Mission-Critical Systems. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999.

ISBN 0-13-020091-3. URL http:

//www.phptr.com/ptrbooks/ptr_
0130200913.html.

Brown:1997:ODA

[Bro97] Lynnwood Brown. Oracle database administration on UNIX systems. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1997. ISBN 0-13-244666-9. xxi + 189 pp. LCCN QA76.9.D3 B7826 1997.

Brown:1999:ODA

[Bro99] Lynnwood Brown. Oracle8 database administration on Windows NT. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-927443-X. 400 pp. LCCN QA76.9.D3B7827 1998. US$39.99. URL http:

//www.phptr.com/ptrbooks/ptr_
013927443x.html.

Burleson:1996:HPO

[Bur96] Donald K. Burleson. Highperformance Oracle database applications. Coriolis Group Books, Scottsdale, AZ, USA, 1996. ISBN 1-57610-100-2. xiii + 461 pp. LCCN QA76.9.D3 B874 1996.

Burleson:1999:OSA

[Bur99] Donald K. Burleson. Oracle SAP Administration. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-696-X. xiii + 197 pp. LCCN QA76.9.D3 B8753 1999. US$32.95. URL http://www.

oreilly.com/catalog/sapadm.

Burleson:2001:UOD

[Bur01] Donald K. Burleson. Unix for Oracle DBAs: pocket reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2001. ISBN 0-596-00066-9. v + 104 pp. LCCN QA76.9.D3 B87 2001.

Corey:1997:ODW

[CA97a] Michael J. Corey and Michael Abbey. Oracle Data Warehousing: A practical guide to successful data warehouse analysis, build, and roll-out. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 1997. ISBN 0-07-882242-4. xxiv + 356 pp. LCCN QA76.9.D3C6735 1997. US$34.95.

URL http://www.osborne.com/
oracle/index.htm.

Corey:1997:OBG

[CA97b] Michael J. Corey and Michael Abbey. Oracle8: A Beginner’s Guide. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 1997. ISBN 0-07-882393-5. xxviii + 767 pp. LCCN QA76.9.D3A24 1997. US$39.99.

URL http://www.osborne.com/
oracle/index.htm.

Corey:1998:ODW

[CAAT98] Michael J. Corey, Michael Abbey, Ian Abramson, and Ben Taub. Oracle8 Data Warehousing. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 1998. ISBN 0-07-882511-3. 686 pp. LCCN QA76.9.D37O73 1998. US$44.99.

URL http://www.osborne.com/
oracle/index.htm.

Corey:1995:TO

[CAD95] Michael J. Corey, Michael Abbey, and Dan J. Dechichio. Tuning Oracle. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 1995. ISBN 0-07-881181-3. xxx + 293 pp. LCCN QA76.9.D3 C6942 1995. US$29.95.

URL http://www.osborne.com/
oracle/index.htm. Covers Oracle

7.x.

Corey:1998:OT

[CADA98] Michael J. Corey, Michael Abbey, Daniel J. Dechichio, and Ian Abramson. Oracle8 Tuning. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 1998. ISBN 0-07-882390-0. xxvi + 608 pp. LCCN QA76.9.D3O72 1998. US$44.99.

URL http://www.osborne.com/
oracle/index.htm.

Carter:1993:PSO

[Car93] John (John R.) Carter. Programming in SQL with Oracle, Ingres, and dBase IV. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1993. ISBN 0-13-014325-1. xi + 308 pp. LCCN QA76.73.S67 C37 1993 Bar.

Clauson:1995:DOI

[CCD^{+}95] Roger Clauson, Joe Cornell, Mark Dreyer, Rick Fisher, John Greco, Mike Kangas, Bill Koch, Tim Marciniak, Ray Sullivan, and Curt Thompson. Detailed Oracle 7 implementation plan. Technical report, Minnesota Dept. of Transportation, Minneapolis, MN, USA, October 13, 1995. 42 pp.

Corrigan:1993:OPT

[CG93] Peter Corrigan and Mark Gurry. ORACLE Performance Tuning. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, September 1993. ISBN 1-56592-048-1. xxxv + 603 pp. LCCN QA76.9.D3 C67 1993. US$34.95.

Chapuis:1996:OUA

[Cha96] Roger Chapuis. Oracle 7: utilisation and administration. Guide de formation. Editions Laser, Saint Dénis, France, 1996. ISBN 2-910525-24-4. 347 pp. LCCN ????

Chang:2000:OXH

[Cha00] Ben Chang. Oracle XML handbook. Osborne/McGraw-Hill, Berkeley, CA, USA, 2000. ISBN 0-07-212489-X. xxii + 357 pp. LCCN QA76.76.H94 O73 2000.

Chu:1995:FDG

[CL95] Ken Chu and Gina Lim. Forms developer’s guide: release 4.5. Oracle Corp., Redwood City, CA, USA, 1995. various pp. Part No. A32505-2.

Chu:1995:FAT

[CLC^{+}95] Ken Chu, Gina Lim, Mark Clark, et al. Forms advanced techniques: release 4.5. Oracle Corp., Redwood City, CA, USA, February 1995. various pp. Part No. A32506-2.

Cobb:1995:EOD

[Cob95] Cobb Group. Exploring Oracle Developer/2000, Designer/2000. Cobb Group, Louisville, KY, USA, 1995. ISSN 1084-1962. various pp.

Cox:1995:OWS

[Cox95] Thomas B. Cox. Oracle workgroup server handbook. Osborne/ McGraw-Hill, Berkeley, CA, USA, 1995. ISBN 0-07-881186-4. xxii + 287 pp. LCCN QA76.9.D3 C693 1995 Bar. Covers Oracle 7 for Windows.

Coppola:1995:GDG

[CP95] Amy Coppola and Andy Page. A guide to Developer/2000 generation: release 1. Oracle Corp., Redwood City, CA, USA, 1995. various pp.

Chang:2001:OXH

[CSK01] Ben Chang, Mark Scardina, and Stefan Kiritzov. Oracle9i XML handbook. Osborne/McGraw-Hill, Berkeley, CA, USA, 2001. ISBN 0-07-213495-X. xxiii + 495 pp. LCCN QA76.76 .H94 C53 2001. CD-ROM includes the official Oracle XML Developer Kit (XDK).

Deleglise:1994:POV

[Del94] Didier Deléglise. Programmer Oracle: versions 6 et 7. Eyrolles, Paris, France, 1994. ISBN 2-212-08813-2. vii + 218 pp. LCCN ????

Duncan:2001:OOS

[DH01] Andy Duncan and Sean Hull. Oracle and Open Source. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel:

+1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, April 2001. ISBN 0-596-00018-9. 406 pp. LCCN QA76.9.D3 D845 2001. US$39.95.

URL http://www.oreilly.com/
catalog/oracleopen.

Dorsey:1997:ODH

[DK97] Paul Dorsey and Peter Koletzke. Oracle Designer/2000 handbook. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 1997. ISBN 0-07-882229-7. xxviii + 610 pp. LCCN QA76.758.D67

1997. URL http://www.osborne.
com/oracle/index.htm.

Donar:2002:TUO

[Don02] Tim Donar. Tru64 UNIX–Oracle9i Cluster quick reference. Compact solutions. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 2002. ISBN 1-55558-272-9 (paperback). xiv + 298 pp. LCCN QA76.76.D63 D66 2002.

Duncan:2002:POD

[DS02] Andrew Duncan and Jared Still. Perl for Oracle DBAs. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00210-6. xvi + 602 pp. LCCN QA76.73.P22 D86 2002 Stacks. US$44.95.

URL  http://safari.oreilly.
com/0596002106;   http://www.
oreilly.com/catalog/oracleperl.

Dye:1999:ODS

[Dye99] Charles Dye. Oracle Distributed Systems. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-432-0. xv + 530 pp. LCCN QA76.73.S67 D59 1999. US$42.95.

URL http://www.oreilly.com/
catalog/ordistsys/.

DIS:1998:OWA

[Dyn98] Dynamic Information Systems, LLC. Oracle Web Application Server Handbook. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 1998. ISBN 0-07-882215-7. xxviii + 684 pp. LCCN QA76.9.D3O734

1998. URL http://www.osborne.
com/oracle/index.htm.

Ensor:1997:OD

[ES97a] Dave Ensor and Ian Stevenson.

Oracle Design. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-268-9. xix + 526 pp. LCCN QA76.9.D26E57 1997. US$39.95. URL http://www.

ora.com/catalog/oracledes/;
http://www.oreilly.com/catalog/
oracledes.

Ensor:1997:ODT

[ES97b] Dave Ensor and Ian Stevenson. Oracle8 Design Tips. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-361-8. xiii + 115 pp. LCCN QA76.9.D26 E575 1997. US$14.95. URL http://www.

ora.com/catalog/oracledt/;
http://www.oreilly.com/catalog/

oracledt. Text edited by Deborah Russell.

Feuerstein:1998:OPS

[FBD98] Steven Feuerstein, John Beresniewicz, and Chip Dawes. Oracle PL/SQL Built-ins Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-456-8. 80 pp. LCCN ???? US$7.95.

URL http://www.oreilly.com/
catalog/orbipr.

Feuerstein:1998:OBP

[FDB98] Steven Feuerstein, Charles Dye, and John Beresniewicz. Oracle Builtin Packages. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, April 1998. ISBN 1-56592-375-8. xxi + 931 pp. LCCN QA76.73.S67 F48 1998. US$46.95. URL http://www.

oreilly.com/catalog/oraclebip.

Covers Oracle8.

Feuerstein:1995:OPS

[Feu95] Steven Feuerstein. Oracle PL/ SQL Programming. Nutshell handbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, September 1995. ISBN 1-56592-142-9. xxvii + 885 pp. LCCN QA76.73.S67 F48 1995.

US$39.95. URL http://www.ora.
com/gnn/bus/ora/item/oraclep.
html.

Feuerstein:1996:AOP

[Feu96] Steven Feuerstein. Advanced Oracle PL/SQL Programming with Packages. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1996. ISBN 1-56592-238-7. xxvi + 661 pp. LCCN QA76.73.P258F48 1996. US$42.95. URL http://www.

ora.com/catalog/advoracle/;
http://www.oreilly.com/catalog/
advoracle.

Feuerstein:1999:OPSb

[Feu99] Steven Feuerstein. Oracle PL/ SQL Programming Guide to Oracle8i Features. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-675-7. xvi + 248 pp. LCCN QA76.73.P258 F484 1999. US$29.95. URL http://www.

oreilly.com/catalog/ornewfeatures.

Feuerstein:2001:OPS

[Feu01] Steven Feuerstein. Oracle PL/SQL Best Practices: Optimizing Oracle Code. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, April 2001. ISBN 0-596-00121-5. xix + 182 pp. LCCN QA76.9.D3 F49 2001. US$19.95.

                                        URL http://www.oreilly.com/
                                        catalog/orbestprac.

Feuerstein:2000:OPS

[FO00] Steven Feuerstein and Andrew Odewahn. Oracle PL/SQL Developer’s Workbook. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 2000. ISBN 1-56592-674-9. xix + 569 pp. LCCN QA76.73.P258 F4875 2000. US$36.95. URL http://www.

                                        oreilly.com/catalog/ordevworkbook.

Feuerstein:1997:OPS

[FP97] Steven Feuerstein and Bill Pribyl. Oracle PL/SQL Programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 1997. ISBN 1-56592-335-9. xxxiii + 987 pp. LCCN QA76.73.S67 F49 1997. US$44.95. URL http://www.

                                        ora.com/catalog/oraclep2/;
                                        http://www.oreilly.com/catalog/
                                        oraclep2.

Feuerstein:2002:OPS

[FP02] Steven Feuerstein and Bill Pribyl.

Oracle PL/SQL Programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, third edition, 2002. ISBN 0-596-00381-1. xxiii + 989 pp. LCCN QA76.73.S67 F49 2002. US$54.95.

URL  http://safari.oreilly.
com/0596003811;   http://www.
oreilly.com/catalog/oraclep3.

Feuerstein:2005:OP

[FP05] Steven Feuerstein and Bill Pribyl.

Oracle PL/SQL Programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, fourth edition, 2005. ISBN 0-596-00977-1 (paperback). 1104 (est.) pp. LCCN ???? URL http://www.

oreilly.com/catalog/oraclep4/.

Feuerstein:1999:OPSa

[FPD99] Steven Feuerstein, Bill Pribyl, and Chip Dawes. Oracle PL/SQL Language Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-457-6. v + 94 pp. LCCN QA76.9.D3 F493 1999. US$7.95.

URL http://www.oreilly.com/
catalog/orbilangpr/.

Finkelstein:1997:OPO

[FSG97] Richard Finkelstein, Kasu Sista, and Rick Greenwald. Oracle Power Objects Developer’s Guide. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 1997. ISBN 0-07-882163-0. xxvii + 428 pp. US$39.95. URL http://www.

                                        osborne.com/oracle/index.htm.

Gurry:1996:OPT

[GC96] Mark Gurry and Peter Corrigan. Oracle Performance Tuning. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, November 1996. ISBN 1-56592-237-9. xxii + 940 pp. LCCN QA76.9.D3G873 1996. US$44.95. URL http://www.

                                        oreilly.com/catalog/oracle2.

Gennick:1999:OSD

[Gen99] Jonathan Gennick. Oracle SQL*Plus: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-578-5. xxi + 502 pp. LCCN QA76.73.S67 G46 1999. US$34.95. URL http://www.

                                        oreilly.com/catalog/orsqplus/.

Gennick:2000:OSP

[Gen00] Jonathan Gennick. Oracle SQL*Plus Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel:

+1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, April 2000. ISBN 1-56592-941-1. iv + 90 pp. LCCN QA76.73.S67 G46 2000. US$9.95.

URL http://www.oreilly.com/
catalog/orsqlpluspr.

Gennick:2002:OSP

[Gen02] Jonathan Gennick. Oracle SQL*Plus Pocket Reference: A Guide to SQL*Plus Syntax. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2002. ISBN 0-596-00441-9. 120 (est.) pp. LCCN ???? US$12.95.

URL  http://safari.oreilly.
com/0596004419;   http://www.
oreilly.com/catalog/orsqlpluspr2.

Greenwald:1997:MOP

[GH97] Rick Greenwald and Robert Hoskin.

Mastering Oracle Power Objects. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1997. ISBN 1-56592-239-5. xx + 485 pp. LCCN QA76.9.D3G728 1997. US$39.95. URL http://www.

ora.com/catalog/oraclepo/;
http://www.oreilly.com/catalog/
oraclepo.

Greenwald:2002:ON

[GK02] Rick Greenwald and David C.

Kreines. Oracle in a Nutshell. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00336-6. 912 (est.) pp. LCCN ???? US$49.95. URL http://www.

                                        oreilly.com/catalog/oracleian.

Gennick:2001:OSD

[GM01] Jonathan Gennick and Sanjay Mishra. Oracle SQL*Loader: The Definitive Guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, April 2001. ISBN 1-56592-948-9. xvii + 251 pp. LCCN QA76.73.S67 G46 2001. US$34.95.

                                        URL http://www.oreilly.com/
                                        catalog/orsqlloader.

Greene:1995:ODS

[Gre95] Joseph B. Greene. Oracle DBA survival guide. SAMS Publishing, Indianapolis, IN 46268, USA, 1995. ISBN 0-672-30681-6. xx + 769 pp. LCCN QA76.9.D3 G7254 1995.

Greenwald:1999:OEO

[GSS99] Rick Greenwald, Robert Stackowiak, and Jonathan Stern. Oracle Essentials: Oracle8 and Oracle8i. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-708-7. 350 pp. US$34.95.

URL http://www.oreilly.com/
catalog/oressentials.

Greenwald:2001:OEO

[GSS01] Rick Greenwald, Robert Stackowiak, and Jonathan Stern. Oracle Essentials: Oracle9 i, Oracle8 i and Oracle8. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, June 2001. ISBN 0-596-00179-7. xvii + 364 pp. LCCN QA76.9.D3 G74 2001. US$34.95.

URL http://www.oreilly.com/
catalog/oressentials2.

Guan:1997:EAL

[Gua97] Weihe Guan. Evaluating ARC-to-Oracle linkages. Geo info systems, 7 (2):38–44, February 1997.

Gurry:2002:OST

[Gur02] Mark Gurry. Oracle SQL Tuning: Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00268-8. vi + 102 pp. LCCN QA76.73.S67 G87 2002 Stacks. US$12.95.

URL  http://safari.oreilly.
com/0596002688;   http://www.
oreilly.com/catalog/orsqltunpr.

Humphrey:1995:GDG

[H^{+}95] Laura Humphrey et al. Graphics 2.5 developer’s guide: release 2.5. Oracle Corp., Redwood City, CA, USA, 1995. various pp. Part No. A32482-2.

Harvey:1998:DGO

[HB98] Dennis Harvey and Steve Beitler. The Developer’s Guide to Oracle Web Application Server 3. Addison-Wesley, Reading, MA, USA, 1998. ISBN 0-201-30960-2. xix + 709 pp. LCCN QA76.76.D47H397 1998. US$44.95.

Hursch:1987:WO

[HH87] Jack L. Hursch and Carolyn J. Hursch. Working with Oracle. TAB Professional and Reference Books, Blue Ridge Summit, PA, USA, 1987. ISBN 0-8306-2916-5. xi + 228 pp. LCCN QA76.9.D3 H868 1987. US$19.60.

Hursch:1989:WOV

[HH89] Jack L. Hursch and Carolyn J. Hursch. Working with Oracle version 6.0. Windcrest/McGraw-Hill, Blue Ridge Summit, PA, USA, 1989. ISBN 0-8306-3246-8. xvi + 384 pp. LCCN QA76.9.D3 H8683 1989. US$21.95.

Hursch:1991:OSD

[HH91] Carolyn J. Hursch and Jack L. Hursch. Oracle SQL developer’s guide. Windcrest/McGraw-Hill, Blue Ridge Summit, PA, USA, 1991. ISBN 0-8306-2566-6 (hardcover), 0-8306-2529-1 (paperback). xv + 240 pp. LCCN QA76.73 .S67 H86 1991. US$32.95 (hardcover), US$21.60 (paperback).

Hipsley:1996:DCS

[Hip96] Paul Hipsley. Developing client/ server applications with Oracle Developer/2000. SAMS Publishing, Indianapolis, IN 46268, USA, 1996. ISBN 0-672-30852-5. xxvi + 640 pp. LCCN QA76.9.C55H57 1996.

Hartley:1992:OSP

[HM92] Tim Hartley and Tim Martyn.

ORACLE/SQL: a professional programmer’s guide. McGraw-Hill, New York, NY, USA, 1992. ISBN 0-07-040775-4 (cloth), 0-07-040767-3 (softcover). xx + 543 pp. LCCN QA76.9.D3 H3652 1992. US$49.95 (cloth), US$34.95 (softcover).

Ingram:1997:HPO

[Ing97] Jonathan Ingram. High performance Oracle database automation. Coriolis Group Books, Scottsdale, AZ, USA, 1997. ISBN 1-57610-152-5. xxii + 419 pp. LCCN QA76.9.D3 I53433 1997.

Jain:1997:PWV

[Jai97] Manoj K. Jain. Paradox 7 for Windows 95 vs. personal Oracle 7 for Windows 95. Thesis (m.s.), Hofstra University, Westport, CT, USA, 1997. iv + 73 pp.

Koch:1995:OCR

[KL95] George Koch and Kevin Loney. Oracle: the complete reference. Osborne/McGraw-Hill, Berkeley, CA, USA, third edition, 1995. ISBN 0-07-882097-9 (paperback). xxxvi + 1066 pp. LCCN QA76.9.D3 K62 1995. US$34.95. Covers versions 7, 7.1, and 7.2.

Koch:1997:OCRa

[KL97a] George Koch and Kevin Loney. Oracle: The Complete Reference: Electronic Edition. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 1997. ISBN 0-07-882285-8. xxxv + 1136 pp. LCCN QA76.9.D3K62 1997. US$54.99. URL http://www.

osborne.com/oracle/index.htm.

Koch:1997:OCRb

[KL97b] George Koch and Kevin Loney. Oracle8: The Complete Reference. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 1997. ISBN 0-07-882396-X. xxxvii + 1300 pp. LCCN QA76.9.D3K6182 1997. US$59.99. URL http://www.

osborne.com/oracle/index.htm.

Kreines:1999:ODA

[KL99] David Kreines and Brian Laskey. Oracle Database Administration: The Essential Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-516-5. xvii + 560 pp. LCCN QA76.9.D3 K7344 1999. US$34.95. URL http://www.

oreilly.com/catalog/oradba/.

Kline:1995:OCD

[Kli95] Kevin E. Kline. Oracle’s cooperative development environment: a reference and user’s guide. Datamation book series. Butterworth-Heinemann, Boston, MA, USA, 1995. ISBN 0-7506-9500-5. xiii + 366 pp. LCCN QA76.76.D47 K564 1995.

Kolste:1995:OPO

[KP95] Bruce Kolste and David Peterson. Oracle Power Objects handbook. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 1995. ISBN 0-07-882089-8. xix + 567 pp. LCCN QA76.9.D3 K647

1995. URL http://www.osborne.
com/oracle/index.htm.

Kreines:2000:OSE

[Kre00] David Kreines. Oracle SQL: The Essential Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, October 2000. ISBN 1-56592-697-8. xxxiii + 382 pp. LCCN QA76.73.S67 K74 2000. US$39.95. URL http://www.

oreilly.com/catalog/orsqlter.

Kreines:2005:ODP

[Kre05] David C. Kreines. Oracle DBA pocket guide. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2005. ISBN 0-596-10049-3. vi + 145 pp. LCCN QA76.9.D3 K744 2004.

Krohn:1993:UOT

[Kro93] Mike Krohn. Using the Oracle toolset. Addison-Wesley, Reading, MA, USA, 1993. ISBN 0-201-56538-2. xii + 327 pp. LCCN QA 76.76 A65 K76 1993. US$40.00.

Kuhn:2002:ORP

[KS02] Darl Kuhn and Scott Schulze. Oracle RMAN: Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00233-5. viii + 111 pp. LCCN QA76.9.B32 K84 2002 Stacks. US$12.95.

URL  http://safari.oreilly.
com/0596002335;   http://www.
oreilly.com/catalog/rman.

Lakshman:2002:OJD

[Lak02] Bulusu Lakshman. Oracle and Java development. SAMS Publishing, Indianapolis, IN 46268, USA, 2002. ISBN 0-672-32117-3. xv + 438 pp. LCCN QA76.9.D3 L35 2002.

Lomasky:1998:OS

[LK98] Brian Lomasky and David C. Kreines. Oracle Scripts. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, May 1998. ISBN 1-56592-438-X. xvi + 185 pp. LCCN QA76.73.S67 L66 1998. US$29.95.

URL http://www.oreilly.com/
catalog/oraclescrp. Covers Or-

acle8.

Lockman:1997:TYO

[Loc97] David Lockman. Teach yourself Oracle8 database development in 21 days. SAMS Publishing, Indianapolis, IN 46268, USA, 1997. ISBN 0-672-31078-3 (paperback). xxviii + 665 pp. LCCN QA76.9.D3 L636 1997. US$49.99, CAN$70.95, UK£46.95.

Loney:1994:ODHb

[Lon94a] Kevin Loney. Oracle DBA handbook: 7.3 edition. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 1994. ISBN 0-07-882289-0. xxii + 720 pp. LCCN QA76.9.D3L663 1994b.

URL http://www.osborne.com/
oracle/index.htm.

Loney:1994:ODHa

[Lon94b] Kevin Loney. Oracle DBA handbook: What every system administrator needs to know for effective and efficient database management. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 1994. ISBN 0-07-881182-1 (hardcover). xxii + 672 pp. LCCN QA76.73.S35 L6 1994. US$34.95. URL http://www.

osborne.com/oracle/index.htm.

Loney:1998:ODH

[Lon98] Kevin Loney. Oracle8 DBA handbook. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 1998. xxii + 822

pp. URL http://www.osborne.
com/oracle/index.htm.

Loney:19xx:ODH

[Lonxx] Kevin Loney. Oracle8 DBA Handbook. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 19xx. ISBN 0-07-882406-0. 822 pp. LCCN ???? US$44.99.

URL http://www.osborne.com/
oracle/index.htm.

Luers:1995:EO

[Lue95] Tom Luers. Essential Oracle 7. SAMS Publishing, Indianapolis, IN 46268, USA, 1995. ISBN 0-672-30873-8. xix + 485 pp. LCCN QA76.9.D3L84 1995.

Mishra:2002:MOS

[MB02] Sanjay Mishra and Alan Beaulieu. Mastering Oracle SQL. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00129-0. xiv + 321 pp. LCCN QA76.73.S67 M57 2002 Stacks. US$34.95.

                                        URL  http://safari.oreilly.
                                        com/0596001290;   http://www.
                                        oreilly.com/catalog/mastorasql.

McCullough:1994:OPS

[McC94] Alastair E. McCullough. Oracle PL/ SQL for SQL*Forms: an introduction to PL/SQL for those using SQL *Forms version 3.0. McGraw-Hill, New York, NY, USA, 1994. ISBN 0-07-707925-6. xx + 169 pp. LCCN QA76.73.S67 M33 1994. UK£22.50.

McClanahan:1996:ODG

[McC96] David McClanahan. Oracle developer’s guide. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 1996. ISBN ???? xviii + 794 pp. LCCN QA76.9.D3 M39375 1996.

McDonald:2003:MOP

[McD03] Connor McDonald. Mastering Oracle PL/SQL: Practical Solutions. Apress, Berkeley, CA, USA, 2003.

ISBN 1-59059-217-4. 616 (est.) pp. US$49.99. URL http://www.

apress.com/book/bookDisplay.
html?bID=276.

Morrison:1998:GOI

[MM98] Joline Morrison and Mike Morrison.

A guide to Oracle: including Developer/2000 and the Oracle Web-Server and DBA utilities. Course Technology, Cambridge, MA, USA, 1998. ISBN 0-7600-4992-0. xvi + 584 pp. LCCN ????

Mahapatra:2000:OPP

[MM00] Tushar Mahapatra and Sanjay Mishra. Oracle Parallel Processing. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2000. ISBN 1-56592-701-X. xvii + 268 pp. LCCN QA76.58 .M34 2000 Sci-Eng. US$32.95.

URL http://www.oreilly.com/
catalog/oraclepp.

McDaniel:2002:TPR

[MM02] Jim McDaniel and Patrick Mc-Grath. TOAD Pocket Reference for Oracle. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00337-4. 128 (est.) pp. LCCN ???? US$12.95. URL http://www.

oreilly.com/catalog/tdpr.

Moltyaner:1997:ODT

[MPK97] Lev Moltyaner, Mark Peco, and Dag Hoftun Knutsen. Oracle Developer/2000: travaux pratiques: cours 236. Support de cours. Learning Tree International, Clichy (Hautsde-Seine), France, 1997. ISBN 2-87672-666-1. various pp. LCCN ????

McDaniel:2005:TPR

[MSMS05] Jim McDaniel, Jeff Smith, Patrick McGrath, and Bert Scalzo. TOAD pocket reference for Oracle. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, second edition, 2005. ISBN 0-596-00971-2 (paperback). vi + 122 pp. LCCN QA76.9.D3 M395 2005.

Muench:2000:BOX

[Mue00] Steve Muench. Building Oracle XML Applications. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, September 2000. ISBN 1-56592-691-9. xvi + 792 pp. LCCN QA76.76.H94 M75 2000. US$44.95. URL http://www.

                                        oreilly.com/catalog/orxmlapp.

Includes CD-ROM.

Muller:1996:ODH

[Mul96] Robert J. Muller. Oracle Developer/2000 handbook. Osborne/ McGraw-Hill, Berkeley, CA, USA, 1996. ISBN 0-07-882180-0. xiv + 446 pp. LCCN QA76.76.D47M82 1996.

Muller:1997:ODH

[Mul97] Robert J. Muller. Oracle Developer/2000 Handbook. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, second edition, 1997. ISBN 0-07-882326-9 (paperback). xvii + 728 pp. LCCN QA76.76.D47M82

1997. URL http://www.osborne.
com/oracle/index.htm.

Odewahn:1999:OWA

[Ode99] Andrew Odewahn. Oracle Web Applications: PL/SQL Developer’s Introduction. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1999. ISBN 1-56592-687-0. 256 pp. US$29.95. URL http://www.

oreilly.com/catalog/oracleweb.

Anonymous:1994:OD

[Ora94] Oracle developer, 1994. ISSN 1078-7518. Pinnacle Publishing, Inc., Kent, WA, USA.

Oracle:1995:PO

[Ora95] Oracle Corporation. Personal Oracle 7. Oracle Corp., Redwood Shores, CA, USA, version 7.1.4.1.OC for Windows edition, 1995.

Oracle:1997:AODc

[Ora97a] Oracle Corporation. Administer the Oracle 7 database, 1997. 3 computer disks updated installation instructions + global license terms.

Oracle:1997:AODa

[Ora97b] Oracle Corporation. Administer the Oracle 7 database II. Oracle Corp., Redwood Shores, CA, USA, version 1.0, v7.8, CBT for Windows edition, 1997. 3 computer disks updated installation instructions + global license terms.

Oracle:1997:AODb

[Ora97c] Oracle Corporation. Administer the Oracle 7 database II. Oracle Corp., Redwood Shores, CA, USA, version 1.0, CBT for Windows edition, 1997. 3 computer disks updated installation instructions + global license terms.

Oracle:1997:D

[Ora97d] Oracle Corporation. Developer/ 2000, 1997. 1 computer optical laser disc installation notes + global license terms + guidelines for building applications + 1 computer optical laser disc; 4–3/4 in. (Personal Oracle lite).

Oracle:1997:DIF

[Ora97e] Oracle Corporation. Developer/ 2000: introduction to forms 4.5, 1997. 4 computer disks workbook/reference guide.

Oracle:1997:DUR

[Ora97f] Oracle Corporation. Developer/ 2000: using reports 2.5, 1997. 4 computer disks workbook/reference guide.

Oracle:1997:OWS

[Ora97g] Oracle Corporation. Oracle 7 workgroup server, 1997. 1 computer laser optical disc notes + 1 trouble– shooting tips card + Oracle fail safe (Version 2.0.5 for MS Windows NT and Windows 95 (1 computer laser optical disc; 4 3/4 in.

+ release notes)) + Support notes (Dec. ’97/Version 2.0.2.0. (1 computer laser optical disc; 4 3/4 in.)).

Owens:1996:BID

[Owe96] Kevin T. Owens. Building intelligent databases with Oracle PL/SQL, triggers, and stored procedure. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-443631-8 (paperback). xiv + 370 pp. LCCN QA76.9.D3 O92 1996.

Palinski:1997:ODC

[Pal97] John Palinski. Oracle database construction kit. Que Corporation, Indianapolis, IN, USA, 1997. ISBN 0-7897-1419-1. xxi + 655 pp. LCCN QA76.9.D3 P338 1997.

Patrick:1999:SF

[Pat99] John Patrick. SQL Fundamentals. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-096016-0. 820 pp. URL http:

//www.phptr.com/ptrbooks/ptr_
013014469X.html. Includes CD-

ROM. Covers Microsoft Access and Oracle.

Papaj:1997:ODW

[PB97] Robert Papaj and Donald Burleson.

Oracle Databases on the Web: Learn to Create Web Pages That Interface with Database Engines. Coriolis Group Books, Scottsdale, AZ, USA, 1997. ISBN 1-57610-099-5. xvi + 572 pp. LCCN QA76.9.D26P37 1997. US$39.99, CAN$55.99. URL

http://www.coriolis.com/cgi-
win/tipsheet.exe/tipsheet?name=

usoracdb. Includes CD-ROM.

Pribyl:2002:LOP

[PF02] Bill Pribyl and Steven Feuerstein. Learning Oracle PL/SQL. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2002. ISBN 0-596-00180-0. xix + 401 pp. LCCN QA76.73.P258 P74 2002 Stacks. US$39.95.

URL  http://safari.oreilly.
com/0596001800;   http://www.
oreilly.com/catalog/learnoracle.

Page:1998:UO

[PH98] William G. Page and Nathan Hughes. Using Oracle8. Que Corporation, Indianapolis, IN, USA, special edition edition, 1998. ISBN 0-7897-1345-4. xix + 929 pp. LCCN QA76.9.D3P334 1998.

Price:2001:JPO

[Pri01] Jason Price. Java Programming with Oracle SQLJ. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, September 2001. ISBN 0-596-00087-1. xiii + 381 pp. LCCN QA76.73.J38 P74 2001. US$39.95. URL http://www.

oreilly.com/catalog/orasqlj.

Russel:2004:ODB

[RC04] Charlie Russel and Robert Cordingley. Oracle DBA Backup and Recovery Quick Reference. The Prentice Hall PTR Oracle series. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 2004. ISBN 0-13-140304-4. xvi + 336 pp. LCCN QA76.9.D3 R868 2004. US$29.99.

RevealNet:2001:ODC

[Rev01] RevealNet. Oracle DBA Checklists Pocket Reference. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, April 2001. ISBN 0-596-00122-3. 80 pp. LCCN ???? US$9.95.

URL http://www.oreilly.com/
catalog/ordbacheck.

Rodgers:1999:ODD

[Rod99] Ulka Rodgers. Oracle: A Database Developer’s Guide. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1999. ISBN 0-13-841420-3. ???? pp. LCCN

???? URL http://www.phptr.com/
ptrbooks/ptr_0138414203.html.

Ravada:1999:OSE

[RS99] S. Ravada and J. Sharma. Oracle8i spatial: Experiences with extensible databases. Lecture Notes in Computer Science, 1651:355–??, 1999. CODEN LNCSD9. ISSN 0302-9743.

Rosenzweig:2004:OPS

[RS04] Benjamin Rosenzweig and Ellen Silvestrova. Oracle PL/SQL by Example. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, third edition, 2004. ISBN 0-13-117261-1. xxiii + 735 pp. LCCN QA76.73.P258 R68 2004. US$49.99.

Singh:1997:ODG

[S^{+}97] Lave Singh et al. Oracle 7.3 developer’s guide. SAMS Publishing, Indianapolis, IN 46268, USA, 1997. ISBN 0-672-22794-0. xxvi + 966 pp. LCCN QA76.9.D3 S5765 1997.

SAS:1993:SAI

[SAS93] SAS Institute. SAS/ACCESS interface to ORACLE: usage and reference, version 6. SAS Institute, SAS Circle, Box 8000, Cary, NC 27512-8000, USA, second edition, 1993. ISBN 1-55544-580-2. xxii + 261 pp. LCCN QA276.4.S21577 1993.

Seligman:1995:GRM

[Sel95] Russ Seligman. Graphics reference manual: release 2.5. Oracle Corp., Redwood City, CA, USA, 1995. various pp. Part No. A32483-1.

Smine:1994:CDA

[Smi94] Hatem Smine. Conçevoir et developper avec Oracle et Case: Oracle versions 6 et 7. Eyrolles, Paris, France, 1994. ISBN 2-212-08829-9. ix + 268 pp. LCCN ????

Sobecki:1997:NSM

[Sob97] Donna Sobecki. NIST space management system using ORACLE and the World Wide Web. Boulder, Co., 1997. ii + 85 + [3] pp.

Stowe:1996:ODH

[Sto96] Michael W. Stowe. Oracle Developer/2000 handbook. P T R Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-227968-1. xii + 292 pp. LCCN QA76.9.D3S757 1996.

Sturner:1995:OUDa

[Stu95a] Günter Sturner. Oracle 7: a user’s and developer’s guide, including release 7.1. Van Nostrand Reinhold Co., New York, NY, USA, 1995. ISBN 0-442-01962-9. xiv + 424 pp. LCCN QA76.9.D3 S82 1995. Translated and adapted by George Shaw and Patricia Shaw.

Sturner:1995:OUDb

[Stu95b] Günter Sturner. Oracle 7: a user’s and developer’s guide, including version 7.1. International Thomson Publishing, London, New York, Boston, Madrid, …, 1995. ISBN 1-85032-118-3. xiv + 424 pp. LCCN QA76.9.D3 S82 1995. Translated and adapted by George Shaw and Patricia Shaw.

Sturner:1997:DIR

[Stu97] Gunther Sturner. Database and information retrieval. Reader. Database and information retrieval. Deakin University, Geelong, Victoria, Australia, 1997. ii + 58 pp.

Toledo:2000:ONC

[TG00] Hugo Toledo and Jonathan Gennick. Oracle Net8 Configuration and Troubleshooting. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, December 2000. ISBN 1-56592-753-2. 408 pp. US$39.95.

URL http://www.oreilly.com/
catalog/net8trouble.

Theriault:1998:OS

[TH98] Marlene Theriault and William Heney. Oracle Security. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 1998. ISBN 1-56592-450-9. 425 pp. LCCN ???? US$34.95. URL http://www.

                                        oreilly.com/catalog/orasec.

Tjahjadi:1997:DIE

[Tja97] Michael Tjahjadi. The design and implementation of an express to Oracle 8 mapping. Thesis (m.s.), Arizona State University, Tempe, AZ, USA, 1997. xii + 225 pp.

Toledo:1996:ON

[Tol96] Hugo Toledo, Jr. Oracle Networking. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 1996. ISBN 0-07-882165-7. xxvii + 365 pp. LCCN QA76.9.D3T63 1996. US$34.95. URL http://www.

                                        osborne.com/oracle/index.htm.

Urman:1996:OPS

[Urm96] Scott Urman. Oracle PL/SQL programming. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 1996. ISBN 0-07-882176-2. xxvi + 564 pp. LCCN QA76.9.D3U76 1996. URL http://

                                        www.osborne.com/oracle/index.

htm. Covers PL/SQL 2.2 and 2.3.

Urman:1997:OPS

[Urm97] Scott Urman. Oracle8 PL/SQL programming: The essential guide for every Oracle programmer: with new information on performance and tuning. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 1997. ISBN 0-07-882305-6. xxxi + 1022 pp. LCCN QA76.9.D3U758 1997.

URL http://www.osborne.com/
oracle/index.htm. Covers releases

7 through 8.

Velpuri:1997:OT

[VA97] Rama Velpuri and Anand Adkoli. Oracle Troubleshooting. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 1997. ISBN 0-07-882388-9. xix + 454 pp. LCCN QA76.9.D3V4463

1997. URL http://www.osborne.
com/oracle/index.htm.

Velpuri:1998:OBR

[VA98] Rama Velpuri and Anand Adkoli. Oracle8 Backup and Recovery Handbook. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 1998. ISBN 0-07-882389-7 (paperback). xxii + 628 pp. LCCN QA76.9.D3 V444

1998. URL http://www.osborne.
com/oracle/index.htm.

Velpuri:1995:OBR

[Vel95] Rama Velpuri. Oracle backup and recovery handbook. Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 1995. ISBN 0-07-882106-1. xv + 384 pp. LCCN QA76.9.D3 V45 1995. URL http://

www.osborne.com/oracle/index.

htm. For all versions of Oracle 7. Velpuri:1997:OBR

[Vel97] Rama Velpuri. Oracle Backup and Recovery Handbook, 7.3 Edition.

Oracle Press (division of Osborne McGraw-Hill), Berkeley, CA, USA, 1997. ISBN 0-07-882323-4. xvii + 422 pp. LCCN QA76.9.D3V446 1997. US$34.99. URL http://www.

                                        osborne.com/oracle/index.htm.

Whalen:1998:TYO

[WD98] Edward Whalen and Steve Adrien DeLuca. Teach yourself Oracle8 in 21 days. SAMS Publishing, Indianapolis, IN 46268, USA, 1998. ISBN 0-672-31159-3 (paperback). xxvi + 604 pp. LCCN QA76.9.D3 W4784 1998. US$39.99.

Wilson:1997:DBG

[Wil97] Mike Wilson. The difference between God and Larry Ellison: inside Oracle Corporation. Morrow, New York, NY, USA, 1997. ISBN 0-688-14925-1. 385 pp. LCCN HD9696.C62 E57648 1997.

Webb:1991:ODS

[WL91] Kenneth Webb and Lori Lafreniere. Oracle distributed systems: a C programmer’s development guide. Windcrest/McGraw-Hill, Blue Ridge Summit, PA, USA, 1991. ISBN 0-8306-7774-7. xvi + 329 pp. LCCN QA76.73 .S67 W43 1991. US$34.95.